OpenStudy (anonymous):

HELP PLEASE!!! Solve: -4(x + 1) - 3 = -3(x - 4) -19 19 2 0

OpenStudy (mathstudent55):

You need to distribute on both sides. Are you familiar with the distributive property?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so its going to be (-4x - 4) -3=(-3x+12) ?

OpenStudy (mathstudent55):

Great. You did it correctly, but after you distribute, you don't need parentheses. -4x - 4 - 3 = -3x + 12

OpenStudy (mathstudent55):

Now you have two numbers, -4 and -3, on the left side, that can be combined together, so combine them together on the left side.

OpenStudy (anonymous):

ohh ok umm soo you do -4x-7=-3x +12?

OpenStudy (mathstudent55):

Great again. Now we want x-terms on the left side and numbers on the right side. Let's deal with the x-terms first. Since on the right you have -3x, you need to add 3x to the right side to get rid of -3x. That is because -3x plus 3x = 0. In an equation, you need to do the same thing to both sides. Now we add 3x to both sides to get rid of -3x on the right side. -4x - 7 = -3x + 12 +3x +3x ------------------------(add) -x -7 = 12

OpenStudy (mathstudent55):

Do you follow the step above?

OpenStudy (anonymous):

ok so now you have to add seven to twelve?

OpenStudy (mathstudent55):

Correct. Now we do the same we did to the -3x to the -7. We add 7 to get rid of the -7 from the left side, but we need to add 7 to both sides.

OpenStudy (mathstudent55):

-x -7 = 12 +7 +7 --------------(add) -x = 19 Now we only have one step left. We want x, but we have -x on the left side. Now we need to multiply or divide -x by -1 to get x. We must do the same to the other side. Let's divide both sides by -1: -x/-1 = 19/-1 x = ??? (What is 19/-1 ?)

OpenStudy (anonymous):

-19?

OpenStudy (mathstudent55):

Correct. x = -19 Now we can check to see if our solution is correct. -4(x + 1) - 3 = -3(x - 4) We replace x with -19, and we see if the statement is true. -4(-19 + 1) - 3 =? -3(-19 - 4) -4(-18) - 3 =? -3(-23) 72 - 3 =? 69 69 = 69 =? is used as a question "is it equal?" Both sides work out to 69, so the two sides are equal, and the solution is indeed x = -19.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (mathstudent55):

you're welcome